# Divide, Conquer and Unite: Hierarchical Style-Recalibrated Prototype Alignment for Federated Medical Image Segmentation

### v1

@@ 1 Introduction
Medical image segmentation is crucial for computer-assisted diagnosis, and most deep learning-based techniques require large amounts of data for effective training (wang2024cardiovascular; zhao2024ultrasound; zhao2024sam). However, sharing such data among hospitals is often restricted by privacy concerns and regulatory constraints. Federated learning has emerged as a promising strategy for multi-site collaborative training of a global model without compromising patient privacy. A foundational method in this domain, FedAvg (FedAvg_AISTATS17), aggregates model parameters from different clients to form a global model, which is then redistributed to each client for further training. Despite these advances, data heterogeneity remains a significant challenge in real-world scenarios (FLSurveyandBenchmarkforGenRobFair_TPAMI24; zhao2024morestyle). Datasets from different medical sites are often non-IID due to variations in scanners and imaging protocols, leading to different feature distributions across clients (jiang2022harmofl; wang2025pfedsam; zhao2024wia; zhuang2025activessf). This diversity can hinder model convergence, reduce generalization capabilities, and result in suboptimal performance.
@@
To address data heterogeneity, researchers use model representations to correct local training and improve generalization, employing mean feature vectors for privacy (MOON_CVPR21; FedProx_MLSys2020). Subsequent work then introduces prototypes (class-specific mean feature vectors) and federated prototype learning, where client prototypes minimize distances to global or same-class prototypes while maximizing separation from different-class ones (FedProto_AAAI22; FPL_CVPR23). Many researchers have also introduced specialized strategies to further improve training in these settings. For example, one method clusters client features to generate multiple prototypes per domain instead of a single average, and refines feature-level distance by elevating it to an $\alpha$ power, addressing unequal learning due to domain diversity (FedPLVM_NeurIPS24). Another study leverages Gaussian Mixture Models to generate prototypes and soft predictions, then weights these outputs based on both quality and quantity (FedGMKD_NeurIPS24).
@@
In federated learning across multiple medical institutions, variations in scanning protocols, staining methods, and patient populations often amplify feature distribution differences (jiang2023fair; chen2024think; li2025expert). Despite recent advances, many existing approaches still depend on a single-layer or simplified representation, making it challenging to fully capture the multi-scale contextual complexities and style variations in medical image segmentation, where distribution shifts can be more pronounced. We group these limitations into two key challenges and highlight them in Figure 1:
@@
Layerwise Style Bias Accumulation: Style discrepancies from different medical protocols accumulate through intermediate layers, degrading cross-domain feature alignment. Many existing methods only normalize style at input level or align final-layer features (FPL_CVPR23), leaving intermediate-layer style biases unaddressed and hindering model generalization.
@@
Incomplete Contextual Representation Learning: Single-layer feature alignment fails to capture the complete semantic hierarchy needed for medical image segmentation. While lower layers encode crucial local details and higher layers capture abstract structures (li2023d), relying solely on final-layer representations leads to incomplete semantic understanding and compromised performance (sung2024contextrast).
@@
Considering these issues, we propose a novel method called FedBCS, specifically designed for feature heterogeneous federated medical image segmentation. For Layerwise Style Bias Accumulation challenge, we propose a Frequency-domain Style Recalibration (FSR) module, specifically designed for constructing domain-invariant prototypes in federated learning. We explore methods for decoupling and recalibrating domain-specific style variations during federation by learning adaptive style parameters from cross-client feature representations. Given that style variations propagate through network layers and affect prototype quality differently, the FSR module modulates feature styles in the frequency domain for robust prototype construction. First, we decompose intermediate features into content and style components via Fourier Transform, where phase spectra preserve semantic content while amplitude spectra encode style information (lee2023decompose; chen2021amplitude). Building on this, we introduce learnable style parameters to adaptively recalibrate the normalized amplitude during federation, enabling optimal style modulation while maintaining semantic consistency. This strategy effectively eliminates domain-specific style biases while ensuring robust prototype construction across different domains.
@@ 2.1 Data Heterogeneous Federated Learning
Federated learning (FL) provides a privacy-preserving framework for collaborative model training without data centralization (konevcny2016federated). The foundational method FedAvg (FedAvg_AISTATS17) iteratively builds a global model by aggregating locally trained parameters. While effective in homogeneous settings, FedAvg often struggles with data heterogeneity, which frequently arises in medical contexts where varying imaging protocols, staining methods, or patient populations lead to significant divergence across institutions (wang2022personalizing). Based on FedAvg, numerous methods have attempted to address the data heterogeneity problem from different perspectives (FLSurveyandBenchmarkforGenRobFair_TPAMI24). Some approaches focus on training a single global model, aiming to refine the global parameters for improved cross-domain performance (FPL_CVPR23; huang2022learn; GGEUR_CVPR25), while others emphasize fairness, ensuring balanced outcomes across heterogeneous clients. Among these, individual/group fairness (ezzeldin2023fairfed), collaboration fairness (jiang2023fair), and performance fairness (chen2024fair; li2021ditto) are three widely studied notions. Meanwhile, personalization methods (FedBN_ICLR21; huang2023generalizable) learn individualized models tailored to each client’s domain, and domain generalization approaches (FedDG_CVPR21; zhang2023federated) seek robust performance on unseen distributions. In this paper, we address the challenge of data heterogeneity across different medical institutions and concentrate on training a single global model that can robustly perform across these diverse domains.
@@ 2.2 Model Representation Alignment in Federated Learning
In Federated Learning (FL), many methods rely on model representations to enhance local training in heterogeneous data settings (wan2024federated; liao2024rethinking; ye2023heterogeneous). However, sharing raw features can pose privacy concerns. To address this, mean feature vectors are used to retain essential information while minimizing data leakage. Mean feature vectors computed from same-class samples are called prototypes. MOON (MOON_CVPR21) and FedProx (FedProx_MLSys2020) utilize contrastive learning to maximize the agreement between local and global model mean feature vectors, thereby aligning local representations with the global model. FedProto (FedProto_AAAI22) enhances the alignment of client-side training by minimizing the distance between client-side class prototypes and global class prototypes. However, by emphasizing single-domain performance under label skew, these approaches often neglect domain shift, thereby limiting their generalization across diverse domains. Thus, many methods adopt prototype-based approaches to address domain skew. For instance, FPL (FPL_CVPR23) employs server-side, cluster-based prototypes combined with contrastive learning to align same-class representations. However, it relies on a single representation for each client, leading to limited representational capacity and causing unequal learning in heterogeneous domains. By contrast, FedContrast-GPA (FedContrast-GPA_MICCAI23) and FedPLVM (FedPLVM_NeurIPS24) cluster features on the client side to produce multiple prototypes per class, mitigating the drawbacks of single-representation strategies. FedGMKD (FedGMKD_NeurIPS24), on the other hand, generates prototypes using Gaussian Mixture Models. Despite their effectiveness, these methods rely solely on final-layer encoder features and overlook the influence of domain-specific style differences on feature representations. In this work, we focus on finer-grained feature alignment and explicitly handling style-induced variations to enhance federated segmentation across diverse medical domains.
@@ 3.1 Preliminaries
Following typical Federated Learning setup (FedAvg_AISTATS17; FedProx_MLSys2020), we assume there are $M$ clients. Each client $m$ holds a private dataset $\{(x_{i},y_{i})\}_{i=1}^{N_{m}}$, where $N_{m}$ denotes the size of client $m$’s local dataset. The objective of federated learning is to minimize the weighted average of local objectives:
@@ 3.3 Context-aware Dual-Level Prototype Alignment
Motivation. While our FSR module effectively addresses style variations, existing prototype-based methods (FedProto_AAAI22; FPL_CVPR23) typically only utilize final-layer features for prototype construction and alignment. This single-layer approach overlooks the rich hierarchical information distributed across network layers, where both local details and global structures are crucial for medical image segmentation. When constructing prototypes for cross-client alignment, this leads to:
@@
Server-side Prototype Alignment. In federated medical image segmentation, different institutions may focus on different anatomical regions or pathological conditions, leading to varied feature distributions. To achieve consistent cross-client alignment while preserving privacy, we introduce a clustering-based alignment strategy. Using FINCH (sarfraz2019efficient), we group similar prototypes from different clients:
@@ 4.1 Experiment Settings
  • Histology nuclei segmentation: We assembled four public datasets (TCIA (hou2020dataset), CRC (awan2017glandular), KIRC (irshad2014crowdsourcing), TNBC (naylor2018segmentation)) for binary segmentation of cell nuclei in histology images.
  • Prostate MRI segmentation: We used a multi-site dataset from (liu2020ms), which comprises six sources from three public datasets (lemaitre2015computer; litjens2014evaluation; nciisbi2013) for binary segmentation of the prostate.
We treat each data source as a separate client. Due to variations in pathological staining and MRI scanning protocols, images from each client differ visually, yet they maintain a consistent label distribution. The example cases in each domain are presented in Fig. 1. Following (wang2023black; liu2020ms), we adopt the same preprocessing methods and data splits, and directly evaluate the results on the test sets. For more dataset details, please refer to Appendix.
Model. For both segmentation tasks, we use commonly used UNet (ronneberger2015u) as the shared model architecture for training.
Comparison Methods. We compare our method against several SOTA federated learning approaches designed to construct a shared global model, including FedAvg (FedAvg_AISTATS17), FedProx (FedProx_MLSys2020), MOON  (MOON_CVPR21), FedProto (FedProto_AAAI22), HarmoFL (jiang2022harmofl), FPL (FPL_CVPR23), FedContrast (FedContrast-GPA_MICCAI23), FedPLVM (FedPLVM_NeurIPS24) and FedUV (FedUV_CVPR24).
@@
Evaluation Metric. Following (wang2023black), we adopt the widely used Dice score as the evaluation metric for both segmentation tasks. Our evaluation focuses solely on the foreground Dice score. To ensure reproducibility, we fix the random seed for all experiments.
@@ 4.2 Diagnostic Analysis
Communication Cost Analysis. We compare our method’s communication overhead with other federated prototype approaches in Tab. 3 (prostate MRI segmentation) and Fig. 3 (a). As shown in Fig. 3(a), our method achieves better performance while maintaining stable convergence. In terms of communication efficiency (Tab. 3), FedProto (FedProto_AAAI22) and FPL (FPL_CVPR23) generate two prototypes by averaging each class’s features, while FedContrast (FedContrast-GPA_MICCAI23) and FedPLVM (FedPLVM_NeurIPS24) require more prototypes due to local clustering. Although our method leverages multi-level features to address incomplete contextual representation, we maintain efficiency by uploading only two prototypes per class. The results demonstrate that our approach achieves superior segmentation performance without substantial communication overhead.

### v2

# Divide, Conquer and Unite: Hierarchical Style-Recalibrated Prototype Alignment for Federated Medical Segmentation

@@ 1 Introduction
Medical image segmentation is crucial for computer-assisted diagnosis, and most deep learning-based techniques require large amounts of data for effective training (Wang et al. 2024b; Zhao et al. 2024d, c). However, sharing such data among hospitals is often restricted by privacy concerns and regulatory constraints. Federated learning has emerged as a promising strategy for multi-site collaborative training of a global model without compromising patient privacy. A foundational method in this domain, FedAvg (McMahan et al. 2017), aggregates model parameters from different clients to form a global model, which is then redistributed to each client for further training. Despite these advances, data heterogeneity remains a significant challenge in real-world scenarios (Huang et al. 2024; Zhao et al. 2024a). Datasets from different medical sites are often non-IID due to variations in scanners and imaging protocols, leading to different feature distributions across clients (Jiang et al. 2022; Wang et al. 2025; Zhao et al. 2024b; Zhuang et al. 2025). This diversity can hinder model convergence, reduce generalization capabilities, and result in suboptimal performance.
@@
To address data heterogeneity, researchers use model representations to correct local training and improve generalization, employing mean feature vectors for privacy (Li et al. 2021a, 2020). Subsequent work then introduces prototypes (class-specific mean feature vectors) and federated prototype learning, where client prototypes minimize distances to global or same-class prototypes while maximizing separation from different-class ones (Tan et al. 2022; Huang et al. 2023b). Many researchers have also introduced specialized strategies to further improve training in these settings. For example, one method clusters client features to generate multiple prototypes per domain instead of a single average, and refines feature-level distance by elevating it to an $\alpha$ power, addressing unequal learning due to domain diversity (Wang et al. 2024a). Another study leverages Gaussian Mixture Models to generate prototypes and soft predictions, then weights these outputs based on both quality and quantity (Zhang et al. 2024).
@@
In federated learning across multiple medical institutions, variations in scanning protocols, staining methods, and patient populations often amplify feature distribution differences (Jiang et al. 2023; Chen et al. 2024a; Li et al. 2025). Despite recent advances, many existing approaches still depend on a single-layer or simplified representation, making it challenging to fully capture the multi-scale contextual complexities and style variations in medical image segmentation, where distribution shifts can be more pronounced. We group these limitations into two key challenges and highlight them in Figure 1:
@@
Layerwise Style Bias Accumulation: Style discrepancies from different medical protocols accumulate through intermediate layers, degrading cross-domain feature alignment. Many existing methods only normalize style at input level or align final-layer features (Huang et al. 2023b), leaving intermediate-layer style biases unaddressed and hindering model generalization.
@@
Incomplete Contextual Representation Learning: Single-layer feature alignment fails to capture the complete semantic hierarchy needed for medical image segmentation. While lower layers encode crucial local details and higher layers capture abstract structures (Li et al. 2023), relying solely on final-layer representations leads to incomplete semantic understanding and compromised performance (Sung et al. 2024).
@@
Considering these issues, we propose a novel method called FedBCS, specifically designed for feature heterogeneous federated medical image segmentation. For Layerwise Style Bias Accumulation challenge, we propose a Frequency-domain Style Recalibration (FSR) module, specifically designed for constructing domain-invariant prototypes in federated learning. We explore methods for decoupling and recalibrating domain-specific style variations during federation by learning adaptive style parameters from cross-client feature representations. Given that style variations propagate through network layers and affect prototype quality differently, the FSR module modulates feature styles in the frequency domain for robust prototype construction. First, we decompose intermediate features into content and style components via Fourier Transform, where phase spectra preserve semantic content while amplitude spectra encode style information (Lee et al. 2023; Chen et al. 2021). Building on this, we introduce learnable style parameters to adaptively recalibrate the normalized amplitude during federation, enabling optimal style modulation while maintaining semantic consistency. This strategy effectively eliminates domain-specific style biases while ensuring robust prototype construction across different domains.
@@ 2.1 Data Heterogeneous Federated Learning
Federated learning (FL) provides a privacy-preserving framework for collaborative model training without data centralization (Konečnỳ et al. 2016). The foundational method FedAvg (McMahan et al. 2017) iteratively builds a global model by aggregating locally trained parameters. While effective in homogeneous settings, FedAvg often struggles with data heterogeneity, which frequently arises in medical contexts where varying imaging protocols, staining methods, or patient populations lead to significant divergence across institutions (Wang et al. 2022). Based on FedAvg, numerous methods have attempted to address the data heterogeneity problem from different perspectives (Huang et al. 2024). Some approaches focus on training a single global model, aiming to refine the global parameters for improved cross-domain performance (Huang et al. 2023b, 2022; Ma et al. 2025), while others emphasize fairness, ensuring balanced outcomes across heterogeneous clients. Among these, individual/group fairness (Ezzeldin et al. 2023), collaboration fairness (Jiang et al. 2023), and performance fairness (Chen et al. 2024b; Li et al. 2021b) are three widely studied notions. Meanwhile, personalization methods (Li et al. 2021c; Huang et al. 2023a) learn individualized models tailored to each client’s domain, and domain generalization approaches (Liu et al. 2021; Zhang et al. 2023) seek robust performance on unseen distributions. In this paper, we address the challenge of data heterogeneity across different medical institutions and concentrate on training a single global model that can robustly perform across these diverse domains.
@@ 2.2 Model Representation Alignment in Federated Learning
In Federated Learning (FL), many methods rely on model representations to enhance local training in heterogeneous data settings (Wan et al. 2024; Liao et al. 2024; Ye et al. 2023). However, sharing raw features can pose privacy concerns. To address this, mean feature vectors are used to retain essential information while minimizing data leakage. Mean feature vectors computed from same-class samples are called prototypes. MOON (Li et al. 2021a) and FedProx (Li et al. 2020) utilize contrastive learning to maximize the agreement between local and global model mean feature vectors, thereby aligning local representations with the global model. FedProto (Tan et al. 2022) enhances the alignment of client-side training by minimizing the distance between client-side class prototypes and global class prototypes. However, by emphasizing single-domain performance under label skew, these approaches often neglect domain shift, thereby limiting their generalization across diverse domains. Thus, many methods adopt prototype-based approaches to address domain skew. For instance, FPL (Huang et al. 2023b) employs server-side, cluster-based prototypes combined with contrastive learning to align same-class representations. However, it relies on a single representation for each client, leading to limited representational capacity and causing unequal learning in heterogeneous domains. By contrast, FedContrast-GPA (Zhou and Zheng 2023) and FedPLVM (Wang et al. 2024a) cluster features on the client side to produce multiple prototypes per class, mitigating the drawbacks of single-representation strategies. FedGMKD (Zhang et al. 2024), on the other hand, generates prototypes using Gaussian Mixture Models. Despite their effectiveness, these methods rely solely on final-layer encoder features and overlook the influence of domain-specific style differences on feature representations. In this work, we focus on finer-grained feature alignment and explicitly handling style-induced variations to enhance federated segmentation across diverse medical domains.
@@ 3.1 Preliminaries
Following typical Federated Learning setup (McMahan et al. 2017; Li et al. 2020), we assume there are $M$ clients. Each client $m$ holds a private dataset $\{(x_{i},y_{i})\}_{i=1}^{N_{m}}$, where $N_{m}$ denotes the size of client $m$’s local dataset. The objective of federated learning is to minimize the weighted average of local objectives:
@@ 3.3 Context-aware Dual-Level Prototype Alignment
Motivation. While our FSR module effectively addresses style variations, existing prototype-based methods (Tan et al. 2022; Huang et al. 2023b) typically only utilize final-layer features for prototype construction and alignment. This single-layer approach overlooks the rich hierarchical information distributed across network layers, where both local details and global structures are crucial for medical image segmentation. When constructing prototypes for cross-client alignment, this leads to:
@@
Server-side Prototype Alignment. In federated medical image segmentation, different institutions may focus on different anatomical regions or pathological conditions, leading to varied feature distributions. To achieve consistent cross-client alignment while preserving privacy, we introduce a clustering-based alignment strategy. Using FINCH (Sarfraz et al. 2019), we group similar prototypes from different clients:
@@ 4.1 Experiment Settings
  • Histology nuclei segmentation: We assembled four public datasets (TCIA (Hou et al. 2020), CRC (Awan et al. 2017), KIRC (Irshad et al. 2014), TNBC (Naylor et al. 2018)) for binary segmentation of cell nuclei in histology images.
  • Prostate MRI segmentation: We used a multi-site dataset from (Liu et al. 2020), which comprises six sources from three public datasets (Lemaître et al. 2015; Litjens et al. 2014; Nicholas et al. 2015) for binary segmentation of the prostate.
We treat each data source as a separate client. Due to variations in pathological staining and MRI scanning protocols, images from each client differ visually, yet they maintain a consistent label distribution. The example cases in each domain are presented in Fig. 1. Following (Wang et al. 2023; Liu et al. 2020), we adopt the same preprocessing methods and data splits, and directly evaluate the results on the test sets. For more dataset details, please refer to Appendix.
Model. For both segmentation tasks, we use commonly used UNet (Ronneberger et al. 2015) as the shared model architecture for training.
Comparison Methods. We compare our method against several SOTA federated learning approaches designed to construct a shared global model, including FedAvg (McMahan et al. 2017), FedProx (Li et al. 2020), MOON  (Li et al. 2021a), FedProto (Tan et al. 2022), HarmoFL (Jiang et al. 2022), FPL (Huang et al. 2023b), FedContrast (Zhou and Zheng 2023), FedPLVM (Wang et al. 2024a) and FedUV (Son et al. 2024).
@@
Evaluation Metric. Following (Wang et al. 2023), we adopt the widely used Dice score as the evaluation metric for both segmentation tasks. Our evaluation focuses solely on the foreground Dice score. To ensure reproducibility, we fix the random seed for all experiments.
@@ 4.2 Diagnostic Analysis
Communication Cost Analysis. We compare our method’s communication overhead with other federated prototype approaches in Tab. 3 (prostate MRI segmentation) and Fig. 3 (a). As shown in Fig. 3(a), our method achieves better performance while maintaining stable convergence. In terms of communication efficiency (Tab. 3), FedProto (Tan et al. 2022) and FPL (Huang et al. 2023b) generate two prototypes by averaging each class’s features, while FedContrast (Zhou and Zheng 2023) and FedPLVM (Wang et al. 2024a) require more prototypes due to local clustering. Although our method leverages multi-level features to address incomplete contextual representation, we maintain efficiency by uploading only two prototypes per class. The results demonstrate that our approach achieves superior segmentation performance without substantial communication overhead.